# Localised Natural Causal Learning Algorithms for Weak Consistency Conditions

### v1

@@ 3.1.1 Realising and Interpreting the V-OUS Condition
With V-OUS being one of our consistency conditions that is somewhat strong, and is implied by faithfulness. Without assuming faithfulness, we discuss cases at which the V-OUS property can still arise, and provide basic interpretations.
@@
    (Composition property) for all disjoint $i,j,k,C$, the following holds: $i\mbox{$\>\perp\perp$ }j\,|\,C\And i\mbox{$\>\perp\perp$ }k\,|\,C\Rightarrow i\mbox{$\>\perp\perp$ }j\cup k\,|\,C$.
@@
    (Conditional exchangability) for all non-collider v-configuration $i\sim k\sim j$ in $G_{0}$, the marginal distribution of $P$ on $\{i,k,j\}$ conditioned on $V\backslash\{i,j,k\}$ is exchangable.
@@
The composition property allows the deduction of joint independence from pairwise independence, and is satisfied by some common distributions such as Gaussians. While the exchangability assumption is commonly made when nodes are indistinguishable from one another, such as in Bayesian theory.
@@ 3.2 Construction of a LoNS algorithm
1:Construct $sk(P)$.
2:Apply the V-OUS and collider-stable orientation rule wrt $P$ to assign v-configurations in $sk(P)$.
3:Solve for a DAG $G(P)$ having skeleton $sk(P)$ and satisfy the assignment of v-configurations. If no solution exists, return error.
@@
Generally, Me-LoNS differs from PC [Spirtes et al., 2000] only in determining whether v-configurations in $sk(P)$ is a collider.
@@
Since the running time of greedy versions of SP based on depth-first search [Solus et al., 2021, Lam et al., 2022] are generally dependent on the depth parameter, it is not obvious whether the running time of Me-LoNS is an improvement.
Note that Me-LoNS outputs a DAG, since in the observational causal learning setting we are interested in the corresponding MEC, we can always convert the DAG into CP-DAG which is a graphical object uniquely representing a MEC, for example via the dag2cpdag function in the causal-learn Python package [Zheng et al., 2023].
@@
If Me-LoNS errors, then there is no DAG that satisfies the conditions in Proposition 5, thus providing some means of error control in making assumptions relating distribution $P$ and true causal graph $G_{0}$.
@@ 4 Simulation and Theoretical Comparisons
We will compare Me-LoNS to some existing constraint-based causal learning algorithms both theoretically and via simulations using the causal-learn package [Zheng et al., 2023] in Python. Me-LoNS is implemented as follows:
@@
    (Step 1) The same skeleton discovery function as PC is used.
@@
    (Step 2) A new orientation function based on the new orientation rule is made.
@@
    (Step 3) The scipy.optimize package is used to solve the DAG search problem.
We will be using mixed linear integer programming with a constant objective to solve Step 3 of Me-LoNS. In addition to the layered network (LN) formulation from Manzour et al. [2021], we introduce additional constraints from Step 2 of Me-LoNS as follows:
@@
where $z_{ij}=1$ if $i\xrightarrow[]{}j$, and $z_{ij}=0$ if $i\xleftarrow[]{}j$, and $\mathcal{C,N}$ the set of v-configurations in $sk(P)$ that are assigned to be colliders and non-colliders respectively by the V-OUS and collider-stable orientation rule wrt $P$ in Step 2 of Me-LoNS.
In each comparison, we will simulate data from a different structural equation model (SEM), with different corresponding causal graph $G_{0}$. In 100 tests of 10000 samples each, we compare the percentage of times the algorithms return the consistent output (output is Markov equivalent to $G_{0}$). Whenever conditional independence testing is needed, the fisherz conditional independence test from the package is used throughout with a significance of 0.05. To test for Markov equivalence of the true causal graph $G_{0}$ and output graph of the algorithm, the mec_check function is used.
@@
Although Me-LoNS is deterministic, due to conditional independence testing being used in simulations, the simulation output is non-deterministic.
@@ 4.1 Comparison to the PC algorithm
If $P$ is V-stable, then the outputs of both PC and Me-LoNS are Markov equivalent. Furthermore there exist distribution $P$ and true causal graph $G_{0}$, such that Me-LoNS is consistent but not PC.
@@
In general, PC outputs a representative of a MEC (CP-DAG). Proposition 6 states that under V-stability, Me-LoNS returns a DAG that is of the MEC represented by the CP-DAG output of PC.
To illustrate Proposition 6, we compare Me-LoNS with the PC algorithm from the package using the definiteMaxP orientation rule. The input distribution $P$ will be induced by the following SEM 1, having all the conditional independencies implied by the Markovian property wrt $G_{0}$ in Figure 2, in addition to $X_{1}\mbox{$\>\perp\perp$ }X_{2}\,|\,X_{3},X_{4}$.
@@
Figure 2: SEM 1 corresponds to DAG $G_{0}$.
PC fails to identify the colliders in Figure 2, as shown in Table 1.
@@ 4.2 Comparison to Sparsest Permutation (SP) algorithm
The Sparsest Markov Representation (SMR) assumption is the sufficient and necessary consistency condition for SP [Raskutti and Uhler, 2018], and it is strictly different to the consistency conditions of Me-LoNS in Theorem 1, as follows:
Example 2 (Me-LoNS and SP are different).
(Conditions in Theorem 1 but not SMR) Consider the graph $G_{0}$ in Figure 3 with $P$ implying the conditional independencies $X_{2}\mbox{$\>\perp\perp$ }X_{3}$ and $X_{1}\mbox{$\>\perp\perp$ }X_{3}$ and $X_{2}\mbox{$\>\perp\perp$ }X_{3}\,|\,X_{1},X_{4}$. Here adjacency faithfulness holds. V-OUS holds since there are no non-colliders to check in $G_{0}$, and modified V-stability of $P$ also holds since V-stability of $P$ holds.
$P$ is Markovian to both $G_{0}$ and $G^{\prime}$ where $G^{\prime}$ differs from $G_{0}$ by flipping the edge $X_{3},X_{4}$. $G_{0}$ and $G^{\prime}$ are both the sparsest Markovian graphs to $P$, but are not Markov equivalent, thus SMR does not hold.
Note that this counter example hinges on the fact that singleton transitivity of $P$ does not hold otherwise we would have $X_{2}\mbox{$\>\perp\perp$ }X_{4}\,|\,X_{1}$ or $X_{3}\mbox{$\>\perp\perp$ }X_{4}\mbox{$\>\perp\perp$ }X_{1}$, violating adjacency faithfulness, thus $P$ cannot be Gaussian.
(SMR but not conditions in Theorem 1) Consider the example in the proof of Theorem 2.4 in Raskutti and Uhler [2018] in which SMR holds but adjacency faithfulness is violated.
To illustrate Example 2, we compare Me-LoNS to the implementation of SP in the package, greedy relaxation of sparsest permutation (GRaSP) Lam et al. [2022]. The input distribution $P$ will be induced by the following SEM 2, having all the conditional independencies in Example 2 in addition to $X_{2}\mbox{$\>\perp\perp$ }X_{3}\,|\,X_{1}$ and $X_{1}\mbox{$\>\perp\perp$ }X_{3}\,|\,X_{2}$.
$\epsilon_{i},\phi_{j}\stackrel{{\scriptstyle\text{i.i.d.}}}{{\sim}}Bern(\tfrac{1}{2}),i=1,..,4,j=1,...,5$
@@ 5 Conclusion and Future Work
    Is a strict generalisation the PC algorithm, and is consistent under strictly different conditions than SP.
@@
    Exponential run time which is comparable to the skeleton building step of SGS.
Hence, Me-LoNS provides another option for an algorithm that is consistent strictly beyond faithfulness, but runs in exponential time which is better than the factorial running time of SP algorithm [Raskutti and Uhler, 2018]. Although there exist speed-ups of the SP algorithm, such as ones based on greedy search like GRaSP used in the Section 4, these algorithms are faster at the cost of stronger consistency conditions [Solus et al., 2021, Lam et al., 2022].
Note that the work done is focused on DAGs, it may be possible to extend the work done to ancestral graphs, which represents causal systems with latent variables, since the notion of ordered upward and downward stabilities are well defined for anterial graphs in general [Sadeghi, 2017].
@@ Appendix A Proofs
Proposition 7 ([Lauritzen, 1996]).
If $P$ is Markovian to $G$, then $P$ is pairwise Markovian to $G$: that is, for every non-adjacent $i,j$, $i\mbox{$\>\perp\perp$ }j\,|\,an_{G}(i,j)$.
Proposition 8 ([Verma and Pearl, 1990]).
DAGs $G_{1}$ and $G_{2}$ are Markov equivalent if and only if:
@@
Proposition 9 ([Ramsey et al., 2006]).
$sk(P)=sk(G)$ if and only if $P$ is adjacency faithful wrt $G$.
@@
Let $i,k,j$ be a collider in DAG $G$. Since $P$ is Markovian to $G$, we have that $i\mbox{$\>\perp\perp$ }j\,|\,an_{G}(i,j)$ by the pairwise Markov property in Proposition 7, and by acyclicity of $G$, $k\not\in an_{G}(i,j)$. ∎
@@
If: Let $P$ be V-OUS and collider-stable wrt $G$. Since $sk(P)=sk(G$), it suffices to show Item 2 that, for v-configurations $i\sim k\sim j$ in $sk(P)=sk(G)$:
@@
Only if: Let G satisfy the V-OUS and collider-stable orientation rules wrt $P$. Since then for v-configurations $i\sim k\sim j$ in $G$: From Item 2, we have:
@@
    If $i\sim k\sim j$ is a collider in $G$, then either:
@@
      If $i\sim k\sim j$ is assigned as a collider, collider-stability wrt $G$ then holds for $i\sim k\sim j$.
@@
      If $i\sim k\sim j$ is unassigned, then for all $C$, such that $k\not\in C$ and $i\mbox{$\>\perp\perp$ }j\,|\,C$, we have that $i\mbox{$\>\perp\perp$ }j\,|\,C\cup k$, collider-stability wrt $G$ then holds for $i\sim k\sim j$.
  2. 2.
    If $i\sim k\sim j$ is a non-collider in $G$ then either:
    1. (a)
      If $i\sim k\sim j$ is assigned as a non-collider, V-OUS wrt $G$ then holds for $i\sim k\sim j$.
    2. (b)
      If $i\sim k\sim j$ is unassigned, then for all $C$, such that $k\not\in C$ and $i\mbox{$\>\perp\perp$ }j\,|\,C$, we have that $i\mbox{$\>\perp\perp$ }j\,|\,C\cup k$, V-OUS wrt $G$ then holds for $i\sim k\sim j$.
Thus $P$ is V-OUS and collider-stable wrt $G$, and $sk(P)=sk(G)$ follows from Item 1. ∎
@@
By Propositions 3 we see that $P$ being modified V-stable is equivalent to: DAGs $G$ to which $P$ satisfies $sk(P)=sk(G)$, and V-OUS and collider-stable wrt have unique v-configurations. This is equivalent to all such DAGs are Markov equivalent by Proposition 8. ∎
@@
Let $G_{1}$ and $G_{2}$ be Markov equivalent DAGs. If $P$ is V-OUS and collider-stable wrt $G_{1}$, then $P$ is V-OUS and collider-stable wrt $G_{2}$.
@@
Since by Proposition 8, we have the that $sk(G_{1})=sk(G_{2})$, the v-configurations in $G_{1}$ and $G_{2}$ coincide, thus since $P$ is the same for both $G_{1}$ and $G_{2}$, $P$ being V-OUS and collider-stable wrt $G_{2}$ follows from that of $G_{1}$. ∎
@@
If: Since $P$ is adjacency faithful wrt $G_{0}$, by Proposition 9 and definition of LoNS $sk(P)=sk(G)=sk(G(P))$. $P$ is also V-OUS and collider-stable wrt both $G_{0}$ and $G(P)$, thus $G_{0}$ is Markov equivalent to $G(P)$ by Remark 4.
@@
    (Adjacency faithfulness) By Proposition 8, and since $G(P)$ is Markov equivalent to $G_{0}$, we have that $sk(P)=sk(G(P))=sk(G_{0})$, and by Proposition 9, adjacency faithfulness follows.
@@
    (V-OUS and collider-stable) $P$ being V-OUS and collider-stable wrt $G(P)$ implies $P$ being V-OUS and collider-stable wrt $G_{0}$ by Proposition 10.
@@
    (Modified V-stability) By contradiction, if $P$ is not modified V-stable, then by Remark 4 the Markov equivalence class of $G_{0}$ is not unique, thus $G_{0}$ and $G(P)$ can be not Markov equivalent.∎
To show Proposition 4, we use the following from Sadeghi [2020]:
Proposition 11 ([Sadeghi, 2020]).
If $P$ is exchangable, then $P$ satisfying composition is equivalent to $P$ satisfying upward stability: for all $i,j,k\in V:i\mbox{$\>\perp\perp$ }j\,|\,C\Rightarrow i\mbox{$\>\perp\perp$ }j\,|\,C\cup k$.
@@
If: Since there exists a DAG $G$ to which $P$ satisfies V-OUS and collider-stability and $sk(P)=sk(G)$, Proposition 3 guarantees that a DAG that satisfies the V-OUS and collider-stable orientation rule wrt $P$ exists, and will be returned by Step-3 of Me-LoNS, again by Proposition 3, $P$ will be adajcency faithful and V-OUS and collider-stable wrt this output.
Only if: If there does not exist a DAG $G$ to which $P$ satisfies V-OUS and collider-stability and $sk(P)=sk(G)$, by Proposition 3, there is no DAG that satisfies the V-OUS and collider-stable orientation rule wrt $P$, thus Step 3 of Me-LoNS errors. ∎
@@
    (Me-LoNS generalises PC under V-stability) Under V-stability of $P$, the V-OUS and collider-stable orientation rule wrt $P$ for assigning colliders becomes the following: for $i\sim k\sim j$ in $sk(P)$:
@@
    (Me-LoNS works but not PC) Consider SEM 1, which gives the graph $G_{0}$ in Figure 2 with the set of conditional independencies being $X_{1}\mbox{$\>\perp\perp$ }X_{2}$ and $X_{1}\mbox{$\>\perp\perp$ }X_{2}\,|\,X_{3},X_{4}$ and $X_{3}\mbox{$\>\perp\perp$ }X_{4}\,|\,X_{1},X_{2}$. Thus we see that the V-OUS-collider-stable orientation rule orients all the v-configurations correctly, but PC orients collider $X_{1}\xrightarrow[]{}X_{3}\xleftarrow{}X_{2}$ as a non-collider.∎

### v2

@@ 3.1.1 Realising and Interpreting the V-OUS Condition
V-OUS is implied by faithfulness. Here, without assuming faithfulness, we discuss cases in which the V-OUS property can still arise, and provide basic interpretations.
@@
    (Composition property). For all disjoint $i,j,k,C$, the following holds: $i\mbox{$\>\perp\perp$ }j\,|\,C\And i\mbox{$\>\perp\perp$ }k\,|\,C\Rightarrow i\mbox{$\>\perp\perp$ }\{j,k\}\,|\,C$.
@@
    (Conditional exchangability). For all non-collider v-configuration $i\sim k\sim j$ in $G_{0}$, the marginal distribution of $P$ on $\{i,j,k\}$ conditioned on $V\backslash\{i,j,k\}$ is exchangable.
@@
The composition property allows the deduction of joint independence from pairwise independence, and is satisfied by some common distributions such as Gaussians. The exchangability assumption is commonly made when nodes are indistinguishable from one another, such as in Bayesian theory.
@@ 3.2 Construction of a LoNS algorithm
1:Construct $\text{sk}(P)$.
2:Apply the V-OUS and collider-stable orientation rule wrt $P$ to assign v-configurations in $\text{sk}(P)$.
3:Solve for a DAG $G(P)$ having skeleton $\text{sk}(P)$ and satisfy the assignment of v-configurations. If no solution exists, return error.
@@
Generally, Me-LoNS differs from PC [const] only in determining whether v-configurations in $\text{\emph{sk}}(P)$ is a collider.
@@
Since the running time of greedy versions of SP based on depth-first search [UhlGSP, lam] are generally dependent on the depth parameter, it is not obvious whether the running time of Me-LoNS is an improvement.
Note that Me-LoNS outputs a DAG, and since in the observational causal learning setting we are interested in the corresponding MEC, we can always convert the DAG into CP-DAG which is a graphical object uniquely representing a MEC, for example via the dag2cpdag function in the causal-learn Python package [python].
@@
The orientation rule of Me-LoNS is similar to conservative PC (CPC) [rams] in the sense that:
  1. 1.
    Both assign non-colliders similarly.
  2. 2.
    Both allow for ambiguous or unassigned v-configurations.
  3. 3.
    Both have a criterion when the consistency condition relating distribution $P$ and true causal graph $G_{0}$ is violated; if Me-LoNS errors, there is no DAG that satisfies the conditions in Proposition 5 (by applying the characterisation in Proposition 3).
However, Me-LoNS relaxes the restricted-faithfulness condition of CPC by orienting colliders differently.
@@ 4 Simulation and Theoretical Comparisons
We will compare the consistency conditions of Me-LoNS to some existing constraint-based causal learning algorithms both theoretically and via simulations using the causal-learn package [python] in Python. Me-LoNS is implemented via the following steps:
@@
    Use the same skeleton discovery function as PC.
@@
    Make a new orientation function based on the new orientation rule.
@@
    Use the scipy.optimize package to solve the DAG search problem.
We will be using mixed linear integer programming with a constant objective to solve Step 3 of Me-LoNS. In addition to the layered network (LN) formulation from shoj, we introduce additional constraints from Step 2 of Me-LoNS as follows:
@@
where $z_{ij}=1$ if $i\xrightarrow[]{}j$, and $z_{ij}=0$ if $i\xleftarrow[]{}j$, and $\mathcal{C,N}$ the set of v-configurations in $\text{sk}(P)$ that are assigned to be colliders and non-colliders respectively by the V-OUS and collider-stable orientation rule wrt $P$ in Step 2 of Me-LoNS.
Within each comparison, we will simulate data from the same structural equation model (SEM), with corresponding causal graph $G_{0}$ to obtain a total of 1,000,000 samples. These samples are then subdivided into 100 test units of 10,000 samples each. From these 100 tests, we compare the percentage of tests the algorithms return the consistent output (output is Markov equivalent to $G_{0}$). Whenever conditional independence testing is needed, the fisherz conditional independence test from the package is used throughout with a significance of 0.05. To test for Markov equivalence of the true causal graph $G_{0}$ and the output graph of the algorithm, the mec_check function is used.
@@
Although Me-LoNS is deterministic, due to conditional independence testing being used in simulations, the simulation output is non-deterministic. The simulations aim to investigate how well the theoretical results (stated purely in conditional independencies) hold up under the randomness of conditional independence testing.
@@ 4.1 Comparison to the PC algorithm
If $P$ is V-stable, then the outputs of both PC and Me-LoNS are Markov equivalent. Furthermore, there exist distribution $P$ and true causal graph $G_{0}$, such that Me-LoNS is consistent but not PC.
@@
In general, PC outputs a representative of a MEC (CP-DAG). Proposition 6 states that under V-stability, Me-LoNS returns a DAG that is of the MEC represented by the CP-DAG output of PC regardless of violations of consistency conditions.
@@
Figure 2: SEM 1 corresponds to the DAG $G_{0}$.
To illustrate Proposition 6, we compare Me-LoNS with the PC algorithm from the package using the definiteMaxP orientation rule which orients only definite colliders and definite non-colliders (thus in this setting, PC coincides with CPC). The input distribution $P$ will be induced by SEM 1, having all the conditional independencies implied by the Markovian property wrt $G_{0}$ in Figure 2, in addition to $X_{1}\mbox{$\>\perp\perp$ }X_{2}\,|\,\{X_{3},X_{4}\}$.
PC fails to identify the colliders in Figure 2, due to violation of orientation faithfulness. This is reflected in Table 1.
@@ 4.2 Comparison to Sparsest Permutation (SP) algorithm
The Sparsest Markov Representation (SMR) assumption is the sufficient and necessary consistency condition for SP [UhlSP], and it is strictly different to the consistency conditions of Me-LoNS in Theorem 1, as the following example shows:
Example 2 (Me-LoNS and SP are different/incomparable).
(SMR, but not conditions in Theorem 1). Consider the example from UhlSP. Let $G_{0}$ be the following:
[…]
and $P$ implies the conditional independencies $X_{1}\mbox{$\>\perp\perp$ }X_{3}\,|\,X_{2}$ and $X_{2}\mbox{$\>\perp\perp$ }X_{4}\,|\,\{X_{1},X_{3}\}$ and $X_{1}\mbox{$\>\perp\perp$ }X_{2}\,|\,X_{4}$. It can be seen that SMR holds, but adjacency faithfulness is violated.
(Conditions in Theorem 1, but not SMR). Consider the graph $G_{0}$ in Figure 3, with $P$ implying the conditional independencies $X_{2}\mbox{$\>\perp\perp$ }X_{3}$ and $X_{1}\mbox{$\>\perp\perp$ }X_{3}$ and $X_{2}\mbox{$\>\perp\perp$ }X_{3}\,|\,\{X_{1},X_{4}\}$. Here, adjacency faithfulness holds. V-OUS holds since there are no non-colliders to check in $G_{0}$, and modified V-stability of $P$ also holds since V-stability of $P$ holds.
$P$ is Markovian to both $G_{0}$ and $G^{\prime}$ where $G^{\prime}$ differs from $G_{0}$ by flipping the edge $X_{2}\xrightarrow{}X_{4}$. $G_{0}$ and $G^{\prime}$ are both sparsest Markovian graphs to $P$, but are not Markov equivalent, thus SMR does not hold.
Note that this counter-example hinges on the fact that singleton transitivity of $P$ does not hold otherwise we would have $X_{2}\mbox{$\>\perp\perp$ }X_{4}\,|\,X_{1}$ or $X_{3}\mbox{$\>\perp\perp$ }X_{4}\mbox{$\>\perp\perp$ }X_{1}$, violating adjacency faithfulness, thus $P$ cannot be Gaussian.
$\epsilon_{i},\phi_{j}$ $\stackrel{{\scriptstyle\text{i.i.d.}}}{{\sim}}\mathrm{Bern}(\tfrac{1}{2}),i=1,\ldots,4,j=1,\ldots,5$
@@
To illustrate the incomparability of Me-LoNS and SP from Example 2, we compare Me-LoNS to the implementation of SP in the package, greedy relaxation of sparsest permutation (GRaSP) lam. The input distribution $P$ will be induced by SEM 2, having all the conditional independencies in Example 2 in addition to $X_{2}\mbox{$\>\perp\perp$ }X_{3}\,|\,X_{1}$ and $X_{1}\mbox{$\>\perp\perp$ }X_{3}\,|\,X_{2}$.
@@ 5 Conclusion and Future Work
    It is a strict generalisation the PC algorithm, and is consistent under strictly different conditions than SP.
@@
    It has exponential run time which is comparable to the skeleton building step of SGS.
Hence, Me-LoNS provides another option for an algorithm that is consistent strictly beyond faithfulness, but runs in exponential time which is better than the factorial running time of SP algorithm [UhlSP]. Although there exist speed-ups of the SP algorithm, such as ones based on greedy search like GRaSP used in the Section 4, these algorithms are faster at the cost of stronger consistency conditions [UhlGSP, lam].
Note that the work done is focused on DAGs, it may be possible to extend the work done to ancestral graphs, which represents causal systems with latent variables, since the notion of ordered upward and downward stabilities are well defined for anterial graphs in general [orderfaith].
@@ Appendix A Proofs
Proposition 7 ([lauritzen1996]).
If $P$ is Markovian to $G$, then $P$ is pairwise Markovian to $G$; that is, for every non-adjacent $i,j$, we have $i\mbox{$\>\perp\perp$ }j\,|\,\text{\emph{an}}_{G}(i,j)$.
Proposition 8 ([pearl]).
The DAGs $G_{1}$ and $G_{2}$ are Markov equivalent if and only if:
@@
Proposition 9 ([rams]).
$\text{\emph{sk}}(P)=\text{\emph{sk}}(G)$ if and only if $P$ is adjacency faithful wrt $G$.
@@
Let $i,k,j$ be a collider in the DAG $G$. Since $P$ is Markovian to $G$, we have that $i\mbox{$\>\perp\perp$ }j\,|\,\mathrm{an}_{G}(i,j)$ by the pairwise Markov property in Proposition 7, and by acyclicity of $G$, we have $k\not\in\mathrm{an}_{G}(i,j)$. ∎
@@
If: Let $P$ be V-OUS and collider-stable wrt $G$. Since $\text{sk}(P)=\text{sk}(G$), it suffices to show Item 2 that, for v-configurations $i\sim k\sim j$ in $\text{sk}(P)=\text{sk}(G)$, we have:
@@
Only if: Let G satisfy the V-OUS and collider-stable orientation rules wrt $P$. Since then for v-configurations $i\sim k\sim j$ in $G$: From Item 2, we have the following breakdown:
@@
    Let $i\sim k\sim j$ is a collider in $G$. It is easy to verify that in both cases, where it is assigned as a collider or remains unassigned, that $i\sim k\sim j$ is collider-stable wrt $G$.
  2. 2.
    Let $i\sim k\sim j$ be a non-collider in $G$.
@@
      If $i\sim k\sim j$ is assigned as a non-collider, then $i\sim k\sim j$ is V-OUS wrt $G$.
@@
      Let $i\sim k\sim j$ be unassigned, and $C\subseteq V\setminus\{i,j,k\}$. If $i\mbox{$\>\perp\perp$ }j\,|\,C$, then we must also have $i\mbox{$\>\perp\perp$ }j\,|\,C\cup\{k\}$, otherwise $i\sim k\sim j$ would have been assigned to be a collider. Hence $i\sim k\sim j$ is V-OUS wrt $G$.
Thus $P$ is V-OUS and collider-stable wrt $G$, and $\text{sk}(P)=\text{sk}(G)$ follows from Item 1. ∎
@@
By Proposition 3 we see that $P$ being modified V-stable is equivalent to: DAGs $G$ to which $P$ satisfies $\text{sk}(P)=\text{sk}(G)$, and V-OUS and collider-stable wrt have unique v-configurations. This is equivalent to all such DAGs are Markov equivalent by Proposition 8. ∎
@@
Let $G_{1}$ and $G_{2}$ be Markov equivalent DAGs, with the same distribution $P$. If $P$ is V-OUS and collider-stable wrt $G_{1}$, then $P$ is V-OUS and collider-stable wrt $G_{2}$.
@@
By Proposition 8, $\text{sk}(G_{1})=\text{sk}(G_{2})$ and the v-configurations and colliders in $G_{1}$ and $G_{2}$ coincide. Since $P$ is the same for both $G_{1}$ and $G_{2}$, $P$ is V-OUS and collider-stable wrt $G_{2}$ by virtue that it is for $G_{1}$. ∎
@@
If: Since $P$ is adjacency faithful wrt $G_{0}$, by Proposition 9 and definition of LoNS, $\text{sk}(P)=\text{sk}(G)=\text{sk}(G(P))$. $P$ is also V-OUS and collider-stable wrt both $G_{0}$ and $G(P)$, thus $G_{0}$ is Markov equivalent to $G(P)$ by Remark 4.
@@
    (Adjacency faithfulness). By Proposition 8, and since $G(P)$ is Markov equivalent to $G_{0}$, we have that $\text{sk}(P)=\text{sk}(G(P))=\text{sk}(G_{0})$, and by Proposition 9, adjacency faithfulness follows.
@@
    (V-OUS and collider-stable). Since $P$ is V-OUS and collider-stable wrt $G(P)$, by Proposition 10, $P$ is V-OUS and collider-stable wrt $G_{0}$.
@@
    (Modified V-stability). Appealing to Remark 4, we note that any DAG satisfying the conditions in the remark is the output of a LoNS algorithm; thus by assumption, these DAGS are Markov equivalent to $G_{0}$. ∎
To show Proposition 4, we use the following from exc:
Proposition 11 ([exc]).
If $P$ is exchangable, then $P$ satisfying composition is equivalent to $P$ satisfying upward stability; that is, for all $i,j,k\in V$, we have $i\mbox{$\>\perp\perp$ }j\,|\,C\Rightarrow i\mbox{$\>\perp\perp$ }j\,|\,C\cup\{k\}$.
@@
If: Since there exists a DAG $G$ to which $P$ satisfies V-OUS and collider-stability and $\text{sk}(P)=\text{sk}(G)$, Proposition 3 guarantees that a DAG that satisfies the V-OUS and collider-stable orientation rule wrt $P$ exists, and will be returned by Step-3 of Me-LoNS, again by Proposition 3, $P$ will be adajcency faithful and V-OUS and collider-stable wrt this output.
Only if: If there does not exist a DAG $G$ to which $P$ satisfies V-OUS and collider-stability and $\text{sk}(P)=\text{sk}(G)$, by Proposition 3, there is no DAG that satisfies the V-OUS and collider-stable orientation rule wrt $P$, thus Step 3 of Me-LoNS errors. ∎
@@
    (Me-LoNS generalises PC under V-stability). Under V-stability of $P$, the V-OUS and collider-stable orientation rule wrt $P$ for assigning colliders becomes the following: for $i\sim k\sim j$ in $\text{sk}(P)$, we have
@@
    (Me-LoNS works but not PC). Consider SEM 1, which gives the graph $G_{0}$ in Figure 2 with the set of conditional independencies being $X_{1}\mbox{$\>\perp\perp$ }X_{2}$ and $X_{1}\mbox{$\>\perp\perp$ }X_{2}\,|\,\{X_{3},X_{4}\}$ and $X_{3}\mbox{$\>\perp\perp$ }X_{4}\,|\,\{X_{1},X_{2}\}$. Thus we see that the V-OUS-collider-stable orientation rule orients all the v-configurations correctly, but PC orients the collider $X_{1}\xrightarrow[]{}X_{3}\xleftarrow{}X_{2}$ as a non-collider.∎
References (Supplementary Material)